sum of present ages of a mother and daughter is 47. After 5 years, mother will be 6 more than 2 times her daughter's age. Find their present ages.


Answer by fractalier(6550)

Call the mother's age now, x. The daughter's age is y.
Thus x + y = 47, and x = 47 - y.
In five years,
x + 5 = 6 + 2(y + 5) or
x + 5 = 6 + 2y + 10
x + 5 = 2y + 16
x = 2y + 11
Substituting in, we get
47 - y = 2y + 11
36 = 3y
y = 12
x = 35